OpenStudy (anonymous):

What is the value of the expression when a = 2? 5a^4

OpenStudy (michele_laino):

here, you have to replace a with 2, so we get this: \[\Large 5{a^4} = 5 \cdot \left( {{2^4}} \right) = ...?\]

OpenStudy (anonymous):

whats 2X2?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

then do 4 X 5 =

OpenStudy (anonymous):

and that ur answer

OpenStudy (michele_laino):

we have to compute 2^4 first @QueenofTamaran

OpenStudy (anonymous):

we did

OpenStudy (anonymous):

thats 2x2

OpenStudy (michele_laino):

I'm sorry, 2^4=2*2*2*2=...?

OpenStudy (anonymous):

oh ya lmao i forgot that....my bad

OpenStudy (michele_laino):

hint: \[\Large {2^4} = 2 \cdot 2 \cdot 2 \cdot 2 = 16\] so, substituting, we can write: \[\Large 5{a^4} = 5 \cdot \left( {{2^4}} \right) = 5 \cdot 16 = ...?\]

OpenStudy (anonymous):

80?